Alex runs 124 meters at a speed of 6.5 m/s and then walks 155 meters at a speed of 3.3 m/s. What was her average speed?
1
Expert's answer
2020-09-01T10:37:47-0400

"v_{avr}=\frac{S_1+S_2}{t_1+t_2}=\frac{S_1+S_2}{S_1/v_1+S_2/v_2}="


"=\frac{124+155}{124/6.5+155/3.3}=4.22(m/s)"